Mum is 28 years older than Anthony. Mum is 4 years younger than dad. The three of them have a total age of 84 years. How old is each person?
Answer by LinnW(1048)

let m = Mum's age
d = Dad's age
a = Anthony's age
m = a + 28
d = m + 4
Substitute (a + 28) for m in d = m + 4
d = (a + 28) + 4
d = a + 32
a + m + d = 84
a + (a + 28) + (a + 32) = 84
3a + 60 = 84
subtract 60 from each side
3a = 24
divide each side by 3
a = 8
Since m = a + 28, m = 8 + 28 = 36
Since d = a + 32, d = 8 + 32 = 40